How many hydrogen molecules are needed for complete catalytic hydrogenation of 0.5 moles of thetricylglycerol which forms glycerol, palmitic acid, oleic acid and linoleic acid by hydrolysis?
1
Expert's answer
2019-05-07T02:11:39-0400


n(C15H31COOH) = 0.5 mol.

n(C17H33COOH) = 0.5 mol.

n(C17H31COOH) = 0.5 mol.


C15H31COOH + H2


   0.5mol        Xmol

C17H33COOH + H2 = C17H35COOH

      1mol         1mol


n1(H2) = X = 0.5 mol.


   0.5mol        Ymol

C17H31COOH + 2H2 = C17H35COOH

      1mol         2mol


n2(H2) = X = 1 mol.

n(H2) = n1(H2) + n2(H2) = 0.5 mol + 1 mol = 1.5 mol.

N(H2) = n(H2) × NA = 1.5 mol × 6.02 × 1023 mol-1 = 9.03 × 1023.